40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 5 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  16887   Fri Jun 3 12:13:58 2022 PacoConfigurationCDSFix RFM channels

[Paco, Yuta]

We tried fixing the issue of LSC_TRY and LSC_TRX channels not working. We first did some investigation, and just like previously reported by Chris, narrowed down the issue to the RFM channels coming from c1iscex/c1iscey.

First attempt : FAIL

In our first attempt, we

  1. Tripped ETMX/ETMY watchdogs, ssh to c1iscex/c1iscey and restart the rtcds models.
  2. Since the last step didn't fix things, we decided to do the same thing on c1lsc, c1sus, c1ioo.
  3. After hard rebooting c1ioo and c1lsc (because they died during the stopping of rtcds models), and not experiencing any timing issues (nice), we still don't fix the issue.

Second attempt: Success

A second attempt just followed Koji's previous fix explained here. Basic difference with our first attempt was a hard reboot of c1iscex/c1iscey in addition to the rtcds model restarting. RFM channels were then clear of errors and we recovered our IR transmission channels in the LSC model.

Attachment 1: SoGreen.png
SoGreen.png
  16886   Thu Jun 2 20:05:37 2022 yutaConfigurationPSLIMC input power recovered to 1W, some alignment works

[Paco, Yuta]

We have increased the output power from the PSL table to 951 mW (it was 96.7 mW).
IMC was recovered including WFS, and both arms are flashing nicely in IR.
We tweaked the alignment of GRX and GRY injection to align them with IR, but it was hard.
Right now IR beams are not centered on TMs. We should center them first.

What we did:
Power increase and IMC recovery
 - Replaced a beam splitter which splits the beam into IMC REFL RF PD path and WFS path from R=98% to R=10% one. Reflection goes to RF PD.
 - Put a R=98% beam splitter back into WFS path.
 - We also tried to put a window in front of IMC REFL camera to recover the arrangement in 40m wiki, but the beam reflected from the window was too weak for us to align. So, we decided not to place a window in front of the camera.
 - Attached photos are the IMC REFL path before and after the work.
 - Measured the PSL output power as Koji did in elog 40m/16672. It was measured to be 96.7+/- 0.5 mW.
 - Rotated the HWP using the Universal Motion Controller (it was not possible for us to do it from the MEDM screen). The position was changed from 73.99 deg to 36.99 deg. Output power was measured to be 951 +/- 1 mW
 - IMC locked without any other changes.
 - Changed C1:IOO-WFS_TRIGGER_THRESH_ON to 5000 (was 500). IMC WFS also worked.
 - After running MC WFS relief script, WFS DC offsets and RF offsets are adjusted following the steps in elog 40m/16835. Below are the results.

C1:IOO-WFS1_SEG1_DC.AOFF => -0.0008882080010759334
C1:IOO-WFS1_SEG2_DC.AOFF => -0.0006527877490346629
C1:IOO-WFS1_SEG3_DC.AOFF => -0.0005847311617496113
C1:IOO-WFS1_SEG4_DC.AOFF => -0.0010395992663688955
C1:IOO-WFS2_SEG1_DC.AOFF => -0.0025944841559976334
C1:IOO-WFS2_SEG2_DC.AOFF => -0.003191715502180159
C1:IOO-WFS2_SEG3_DC.AOFF => -0.0036688060499727726
C1:IOO-WFS2_SEG4_DC.AOFF => -0.004011172490815322


IOO-WFS1_I1         :  +1977.7 ->    +2250 (Significant change)
IOO-WFS1_I2         :  +3785.8 ->  +3973.2
IOO-WFS1_I3         :  +2014.2 ->  +2277.7 (Significant change)
IOO-WFS1_I4         :  -208.83 ->  +430.96 (Significant change)
IOO-WFS1_Q1         :  +2379.5 ->  +1517.4 (Significant change)
IOO-WFS1_Q2         :  +2260.4 ->  +2172.6
IOO-WFS1_Q3         :  +588.86 ->  +978.98 (Significant change)
IOO-WFS1_Q4         :  +1654.8 ->  +195.38 (Significant change)
IOO-WFS2_I1         :  -1619.9 ->  -534.25 (Significant change)
IOO-WFS2_I2         :  +1610.4 ->  +1619.8
IOO-WFS2_I3         :  +1919.6 ->  +2179.8 (Significant change)
IOO-WFS2_I4         :    +1557 ->  +1426.6
IOO-WFS2_Q1         :   -62.58 ->  +345.56 (Significant change)
IOO-WFS2_Q2         :  +777.01 ->  +805.41
IOO-WFS2_Q3         :  -6183.6 ->  -5365.8 (Significant change)
IOO-WFS2_Q4         :  +4457.2 ->  +4397.


IFO Alignment
 - Aligned both arms using IR. Both arm flashes at the following, which is consistent with the power increase.
 C1:SUS-ETMX_TRX_OUT_DQ ~1.1
 C1:SUS-ETMY_TRY_OUT_DQ ~0.6
 - With this, we tried to tweak GRX and GRY injection. The following is after the work. We could increase GTRX to 0.204 when the Xarm is aligned to green. This suggests that GRX injection is not aligned nicely yet. But the beams are also not centered on TMs. We should center them first.
 C1:ALS-TRX_OUT_DQ ~0.13
 C1:ALS-TRY_OUT_DQ ~0.07
 - GTRX and GTRY cameras are adjusted to have nicer images. In GRX path, the second and last lens before the PD and CCD was pulled ~ 1 cm behind its original position and both beams realigned. Then, on GRY path, the beam was re-centered on the first and only lens, the whole assembly pushed forward by ~ 2 cm and the beams re-centered.

Next:
 - Center the IR beam on TMs (first by our eyeballs; better to use A2L after arm locking is recovered and coils are balanced)
 - Tweak GRX and GRY injection (restore GRY PZTs?)
 - Install ETMXT camera (if it is easy)
 - Lock Xarm and Yarm (C1:LSC-TRX/Y_OUT needs to be fixed for triggering. Can we use other PDs for triggering?)
 - MICH locking (REFL and AS PDs might need to be re-aligned; they are not receiving much light)
 - RTS model for BHD needs to be updated

Attachment 1: Before.JPG
Before.JPG
Attachment 2: After.JPG
After.JPG
  16885   Wed Jun 1 12:56:44 2022 PacoSummaryElectronicsSTEMlab 125 handout

[Paco, Deeksha]

Yesterday I handed Deeksha a red pitaya (stemlab 125 - 10) to begin her summer work in the lab. The short term goal (~1 week) is to get it to work as a network analyzer and perhaps characterize its ADC/DAC noise spectra.

  16884   Wed Jun 1 11:56:28 2022 yutaUpdateALSShutter driver for GRY replaced

[JC, Yuta]

We replaced a shutter driver for GRY since it stopped working this morning.
We replaced it with a free driver which was sitting on the ITMY table.
The reverse polarity issue of C1:AUX-GREEN_Y_Shutter was fixed by switching one of the switches of the driver from N.O. to N.C.

Also, "Toggle" button was added to IFO_ALIGN.adl so that we can toggle shutters easily to find TEM00. It runs /home/controls/Git/40m/scripts/ALS/ShutterToggler.py.
 

Quote:
 

The green Y shutter now works but in reverese, meaning that sending 1 to C1:AUX-GREEN_Y_Shutter closes the shutter and vice versa. This needs to be fixed.

 

Attachment 1: Screenshot_2022-06-01_12-01-53.png
Screenshot_2022-06-01_12-01-53.png
  16883   Tue May 31 17:30:01 2022 YehonathanUpdateBHDGreen shutters fixed

[Paco, Yehonathan]

We fixed the slow control over the green beam shutters.

At the Y arm the wrong BNC was connected to the shutter driver. We connected the correct BNC to the driver and switched the remote mode. The green Y shutter now works but in reverese, meaning that sending 1 to C1:AUX-GREEN_Y_Shutter closes the shutter and vice versa. This needs to be fixed.

At the X end the problem was a bit more complicated. Previously, the shutter was controlled by c1auxey. We figured that c1auxex has a lot of spare bio channels. We found an Acromag BNC front panel (with wires already soldered to the BNCs) lying around in the lab and installed it on the c1auxex Acromag chassie. We then connected the topmost BNC to channel 0 on XT1111A in the chassie. The BNC was connected to the green shutter driver on the X end.

EPIC channel was added to the c1auxex db file while it was commented out on the psl shutters db file. Modbus was restarted on c1auxex and c1psl. c1psl had to be burt restored to regain MC lock. Now the green X shutter works properly.

  16882   Tue May 31 14:44:02 2022 JCUpdateElectronicsCRT TV / Monitor 6

[Paco, JC]

Paco and I fixed the ethernet cable which was hanging. We stopped models c1x07 and c1su2, realigned the cable to follow the shelf from top, and returned to turn on the computers.

 

Note: There was not a long enough ethernet cable, so we used a female to female adapter and attached 2 ethernet cables.

Quote:

[Yehonathan, Paco, Yuta, JC]

As we were cleaning up this morning, we heard a high pitch sound that turned into a buzz. After searching for where the sound came from, we noticed the CRT TV went out. We swapped this out with a moniter and used a BNC to VGA adapter to display the cameras.

 

  16881   Fri May 27 17:46:48 2022 PacoSummaryComputersCDS upgrade visit, downfall and rise of c1lsc models

[Paco, Anchal-remote, Yuta, JC]

Sometime around noon today, right after cds upgrade planning tour, c1lsc FE fell. We though this was ok because anyways c1sus was still up, but somehow the IFO alignment was compromised (this is in fact how we first noticed this loss). Yuta couldn't see REFL on the camera, and neither on the AP table (!!) so somehow either/all of TT1, TT2, PRM got affected by this model stopping. We even tried kicking PRM slightly to try and see if the beam was nearby with no success.

We decided to restart the models. To do this we first ssh into c1lsc, c1ioo and c1sus and stop all models. During this step, c1ioo and c1sus dropped their connection and so we had to physically restart them. We then noticed DC 0x4000 error in c1x04 (c1lsc iop) and after checking the gpstimes were different by 1 second. We then did stopped the model again, and from fb1 restart all daqd_* services and modprobe -r gpstime, modprobe gpstime, restart c1lsc and start the c1x04 model. This fixed the issue, so we finished restarting all FE models and burt restore all the relevant snap files to today 02:19 AM PDT.

This made the IFO recover its nominal alignment, minus the usual drift.

* The OAF model failed to start but we left it like so for now.

  16880   Fri May 27 17:45:53 2022 yutaConfigurationBHDBHD camera installed, GRY aligned

[JC, Paco, Yuta]

After the IFO recovery (elog 40m/16881), we installed an analog camera for BHD fringe using a BNC cable for old SRMF camera so that we can see it from the control room.
We also aligned AS-LO using LO1,LO2 and AS4.
We then aligned GRY injection to get maximum GTRY.

Maximum TEM00s right now are
 C1:SUS-ETMX_TRX_OUT_DQ ~0.1
 C1:SUS-ETMY_TRY_OUT_DQ ~0.05
 C1:ALS-TRX_OUT_DQ ~0.20
 C1:ALS-TRY_OUT_DQ ~0.18

  16879   Fri May 27 15:53:17 2022 TommyUpdateElectronicsRFSoC MTS Work

With some help from the forums, we printed the status of the DAC MTS sync and were able to determined that our board's vivado design does not have MTS enabled on each tile. To fix this, we will need to construct a new Vivado desgin for the board. We were also warned to "make sure to generate correctly a PL_clock and a PL_sysref with your on board clock synthesizers and to capture them in the logic according to the requirements in PG269" of the RF Manual. From this we should be able to sync the DAC and ADC tiles as desired.

Quote:

We followed the manual's guide for setting up MTS to sync on external signal. In the xrfdc package, we update the RFdc class to have RunMTS, SysRefEnable, and SysRefDisable functions as prescribed on page 180 of the manual. Then, we attempted to run the new functions in the notebook and read the DAC signal outputs on an oscilloscope. The DACs were not synced. We were also unable to get FIFOlatency readings. 

 

  16878   Fri May 27 12:15:30 2022 JCUpdateElectronicsCRT TV / Monitor 6

[Yehonathan, Paco, Yuta, JC]

As we were cleaning up this morning, we heard a high pitch sound that turned into a buzz. After searching for where the sound came from, we noticed the CRT TV went out. We swapped this out with a moniter and used a BNC to VGA adapter to display the cameras.

  16877   Thu May 26 19:55:43 2022 yutaConfigurationBHDOplevs centered, BHD DCPDs are now online

[Paco, Yuta]

We have aligned the IFO (except for LO-AS and GRY), and centered all the oplevs.
We have also restored Gautam's in-air BHD DCPD setup and placed it to ITMY table.
BHD DC PD signals are now online at C1:XO4-MADC1_EPICS_CH4 and CH5. 

Oplevs:
 Aligned the IFO following the steps in elog 40m/16875.
 When we were woking on BHD DCPDs, we lost REFL beam on camera and both arms flashing. Alignment was restored mostly with TT2 pitch.
 We centered all the oplevs after the recovery (see attached).

BHD DCPDs:
 1. We removed a circuit box with M2 ISS photodetector readout board from AP table, in-air BHD photodiodes from optics graveyard. (see LIGO-E2000436 and elog 40m/15493 for wiring diagram)
 2. Taken out temporary two Thorlabs PDA100A used for aligning LO-AS during the vent from ITMY table, and placed the BHD setup in ITMY table (see attached and attached).
 3. DB9 cable (15ft+10ft) was connected from M2 ISS box to anti-aliasing chassis for ADC1 of C1X04 at 1Y2 rack (see attached).
 4. +/-18V power for M2 ISS box was supplied from 1Y1 rack.
 5. BHD DCPD signals are now available at C1:XO4-MADC1_EPICS_CH4 and CH5 (see attached).

Next:
 - Tweak alignment of green Y input to follow Yarm
 - Do LO-AS alignment
 - Centering of PDs everywhere with IFO aligned
 - Update RTS model for BHD

Attachment 1: elog_1Y2.JPG
elog_1Y2.JPG
Attachment 2: elog_BHD.JPG
elog_BHD.JPG
Attachment 3: elog_box.JPG
elog_box.JPG
Attachment 4: Screenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.png
Screenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.png
Attachment 5: Screenshot_2022-05-26_20-35-02.png
Screenshot_2022-05-26_20-35-02.png
  16876   Thu May 26 15:55:10 2022 TommyUpdateElectronicsRFSoC Power Spectrum

Finished building power spectrum analyzer for the RFSoC. There are two things that I would like to address down the road. First is that there is an oscillation between positive and negative voltages at the ADC sampling frequency. This creates an undesirable frequency component at the sampling rate. I have not yet figured out the cause of this positive to negative oscillation and have simply removed half of the samples in order to recover the frequency. Therefore, I would like to figure out the root of this oscillation and remove it. Also, we have a decimation factor of 2 as default by the board which we would like to remove but have been unable to do so.

Example: 8 MHz Square Wave from SRL signal generator.

Attachment 1: Screen_Shot_2022-06-02_at_2.51.23_PM.png
Screen_Shot_2022-06-02_at_2.51.23_PM.png
  16875   Wed May 25 17:34:47 2022 yutaConfigurationBHDIFO recovery - IFO alignment

IFO aligned to maximize flashings, except for GRY and LO-AS.

What we did:
 0. After recovering IMC, C1:IOO-MC_TRANS_SUM was ~1300 with C1:IOO-MC_RFPD_DCMON of ~0.11 (~10% better than what we had during vent). Xarm and Yarm was already flashing and could see the beam at AS and POP cameras.
 1. Aligned ETMX and ITMX to green X input beam to maximize C1:ALS-TRX_OUT, to ~0.19.
 2. Aligned TT2-PR3 to get C1:SUS-ETMX_TRX_OUT flashing at 0.09 at max
 3. Aligned ITMY to have nice POP blinking of MICH at POP camera
 4. Aligned ETMY-PR3 to have C1:SUS-ETMX_TRX_OUT flashing at 0.06 at max
 5. Misaligned ITMY (with +2 in C1:SUS-ITMY_PIT_COMM), and aligned PRM to have PRX (PRM-ITMX cavity) flashing at C1:LSC-ASDC_IN1 at ~20 (offset -70) at max
 6. Misaligned PRM, and aligned SRM to have SRX (SRM-ITMX cavity) flashing at C1:LSC-ASDC_IN1 at ~20 (offset -70) at max
 7. Restored all the alignment. ITMY didn't quite come back, so I need to tweak the alignement to maximize TRY flashing.

Result:
Current alignment is as attached. IR beam at AS, REFL, MCR and green beam at GTRX cameras all seem slightly to the left from monitors, but looks as it was before the pump down.yes GTRY is still clipped, but green Y locks stably. Oplevs were not so useful to recover the alignment. ETMX/Y oplevs did not drifted too much probably because we don't have in-vac steering mirrors.

Next:
 - Tweak alignment of green Y input to follow Yarm
 - Do LO-AS alignment
 - REFL DC is not receiving beam. Re-alignment necessary
 - Oplev centering
 - BHD PDs need to be replaced to lower gain PDs and need to be connected to CDS

Attachment 1: Screenshot_2022-05-25_17-47-57.png
Screenshot_2022-05-25_17-47-57.png
  16874   Wed May 25 16:56:44 2022 PacoConfigurationBHDIFO recovery - IMC alignment

[Yuta, Paco]

We aligned IMC to recover the IFO progressively. First step was to center the MC REFL beamspot on the camera as well as the WFS DC. Then slide MC2 and MC3 together. Below are the alignment slider positions before/after.

  MC1 (before --> after) MC2 (before --> after) MC3 (before --> after)
PIT -0.3398 --> -0.4768 4.1217 --> 4.0737 -1.9808 --> -1.9308
YAW -0.8947 --> -0.7557 -1.2350 --> -1.3350 1.5598 --> 1.5638
  16873   Wed May 25 16:38:27 2022 yutaUpdateSUSSuspensions quick health check

[JC, Yuta]

We did a quick health check of suspesions after the pump down.

Summary:
 - ITMX LRSEN is too bright (~761) and not responding to any optic motions (we knew this before the pump down)
 - ITMY ULCOIL is not working
 - LO1 LLCOIL is not working
 - Damping loops need to be retuned, especially for ETMY (too much damping), SRM, PR3 and AS4 (damping too weak)
 - MC1 sensor outputs are minus instead of plus
 - LO2 OSEMs got stuck during the pump down, but now it is free after some kicks. OSEM sensorr values almost came back (see attached)

What we did:
 1. Kicked optics with C1:SUS-{optic}_{UL,LL,UR,LR,SD}COIL_OFFSET one by one with offsets of +/- 10000 (or 100000), and checked if C1:SUS-{optic}_{UL,LL,UR,LR,SD}SEN_OUT16 move in both directions.

 2. Check if the optic damps nicely.

 3. Attached photo of the note is the result.

Attachment 1: Screenshot_2022-05-25_16-46-13.png
Screenshot_2022-05-25_16-46-13.png
Attachment 2: OSEMcheck.JPG
OSEMcheck.JPG
  16872   Tue May 24 15:21:13 2022 AnchalUpdateBHDFreeswing tests of new SOS started

I modified the script freeSwing.py to use damping loop output switches to free the optic instead of watchdog or coil output filters. This ensures that the free swing test is being done at the nominal position of the optic. I started tests for LO1, LO2, As2, As4, PR2, PR3, and SR2 in a tmux session names freeSwing on rossa.


Note: LO2 face OSEMs are hardly sensitive to any motion right now due to excessive pitch offset required for LO beam. We should relieve this offset to LO1 and rerun this test later.

  16871   Tue May 24 11:04:53 2022 JCUpdateVACBeginning Pumpdown

[JC, Jordan, Paco, Chub]

We began with the pumpdown this morning. We started with the annulus volume and proceeded by using the following:

1. Isolate the RGA Volume by closing of valves VM3 and V7.

2. Opened valves VASE, VASV, VABSSCT, VABS, VABSSCO, VAEV, and VAEE, in that order.

3. Open VA6 to allow P3, FRG3, and PAN to equalize.

4. Turn on RP1 and RP3, rough out annulus volume, once <1 torr turn on TP3. Close V6. Open V5 to pump the annulus volume with TP3.

5. Re route pumping from RP1 and RP3 to the main volume by opening V3 and slowly opening RV1.

6. After ~3.5 hours the pressure in the arms was <500mtorr on both FRG1 and P1a. Turn on TP1 and wait to reach full speed 560 Hz

7. Open V1 with RV2 barely open. The pressure diff between P1a and P2/FRG2 needs to be below 1 torr. This took a couple attempts with the manual valve in different positions. The interlocks were tripped for this reason. Repeat step 7 until the manual gate valve was in a position that throttled pumping enough to maintain the <1 torr differential.

8. Slowly open the manual gate valve over the course of ~ 1 hour. Once the manual gate valve fully opened, pressure in the arms was <1mtorr.

9. V7 was closed, leaving only TP2 to back TP1, while TP3 was used to continue pumping the annuli. Left in that configuration overnight (see attached)

 

We did have to replace gauge PAN becuase it was reading a signal error. In addition, we found the cable is a bit sketchy and has a sharp bend. The signal comes in and out when the cable is fiddled with.

Attachment 1: PUMPDOWN-2022-05-24_16-57-59.png
PUMPDOWN-2022-05-24_16-57-59.png
Attachment 2: C1VAC_Screenshot_2022-05-24_16-59-27.png
C1VAC_Screenshot_2022-05-24_16-59-27.png
  16870   Tue May 24 10:37:09 2022 TegaUpdateVACadded FRG channels to slow channel ini file

[Vacuum gauge sensors]

Paco informed me that the FRG sensor EPICS channels are not available on dataviewer, so I added them to slow channels ini file (/opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini). I also commented out the old CC1, CC2, CC3 and CC4 gauges. A service restart is required for them to become available but this cannot be done right now because it would adversely affect the progress of the upgrade work. So this would be done at a later date.

Quote:

git repo - https://git.ligo.org/40m/vac

Finally incorporated the FRGs into the main modbusIOC service and everything seems to be working fine. I have also removed the old sensors (CC1,CC2,CC3,CC4,PTP1,IG1) from the serial client list and their corresponding EPICS channels. Furthermore, the interlock service python script has been updated so that all occurrence of old sensors (turns out to be only CC1) were replaced by their corresponding new FRG sensor (FRG1) and a redundnacy was also enacted for P1a where the interlock condition is replicated with P1a being replaced with FRG1 because they both sense the main volume pressure.

 

  16869   Mon May 23 13:16:59 2022 PacoUpdateBHDEnd of vent - checks

[Paco, Yuta]

Prep for closing and pump down.

  • Aligned IFO to maximize DC levels.
    • YARM (flashing peak 0.05 with PRM misaligned), XARM (flashing peak 0.06 with PRM misaligned), PRC (PRY flashing -30 @ POPDC, offset -70 and REFL DC 270), SRC (SRY flashing -30 @ POPDC, offset -70), BHD.
  • GTRY clipping
    • We tried moving the alignment of PR3, PR2, ETMY, ITMY to reduce clipping and retain IR flashing. We found it kind of difficult, so we only used the unclipped GTRY temporarily to improve the input YAUX injection after which the YAUX locked. We then restored the clipping in favor of the IR beam alignment.
  • PR3 position
    • PR3 seems to be +1 inch away towards East, nominally placed along North-South, and offset in YAW.
  • Aligned OPLEVs to center at around Mon May 23 13:20:32 2022
  • Snapshot of all cameras in the control room around Mon May 23 13:24:51 2022

[Chub, JC, Jordan, Yuta, Yehonathan, Paco]

Closed in the following order:

  • IMC chamber
  • OMC chamber
  • BS chamber
  • ITMY chamber
  • ITMX chamber

[Yuta, Paco]

After closing the heavy doors, we tried to have GTRY less clipped using PR2, PR3, ITMY and ETMY. During this adventure, we also aligned GRY injection beam by hand. Rotating a waveplate for GRY injection made GRY locking stably at GTRY of ~0.3.

Attachment 1: OSEMs20220523.txt
2022-05-23 13:15:59.531109
 =====BS=====
C1:SUS-BS_ULSEN_OUT16 =   600 
C1:SUS-BS_LLSEN_OUT16 =   575 
C1:SUS-BS_URSEN_OUT16 =   600 
C1:SUS-BS_LRSEN_OUT16 =   635 
C1:SUS-BS_SDSEN_OUT16 =   670 
 =====ITMX=====
C1:SUS-ITMX_ULSEN_OUT16 =   404 
C1:SUS-ITMX_LLSEN_OUT16 =   608 
... 76 more lines ...
Attachment 2: Screenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.png
Screenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.png
Attachment 3: PR3position.JPG
PR3position.JPG
  16868   Fri May 20 20:03:48 2022 PacoUpdateBHDITMY chamber work finished - LO and AS overlapped

[Paco, Anchal, Yuta]

Today, in short we:

  • Recovered alignment of arm cavities, PRC (only ITMX aligned), and then altogether with SRM and PRM aligned to maximize all DCPD levels (AS, POP, REFL, TRX, TRY), but SRC was not flashing and the SRM yaw alignment slider was around its max value, so after recording beam positions on cameras Anchal went into the BS chamber and helped steer the SRC alignment using a combination of SRM, SR2 and AS1. After this every beam was nominally aligned except for LO and AS, which remained to be mode matched.
  • Mode matched LO3-LO4 by hand -- cheeky -- from the ITMY chamber, the final separation between these two mirrors grew by almost 3 inches with respect to the design (!!!) but the LO and AS beams came out nicely. The canonical path used for the steering was LO path, and then we overlapped the beams with the help of a gige basler camera and a couple of DCPDs (Thorlabs).
  • Yuta and Paco started running final checks in preparation for Monday (pumpdown). We aligned the IFO, but noted that using Restore/Misalign sometimes results in hysteresis.. so it is not very reliable for fine alignment modes. Then we optimized DC levels, centered all oplevs, and tweaked Green input alignment on XARM and YARM. The XARM was maximized, but in YARM we could still not get high TEM-00 flashing ...
    • Unfortunately, we discovered a slight clipping of the GTRY beam through PR3 which could mean the current alignment (pointing) is not hitting PR3 center optimally.
  • Attached are the screenshot of current aligned state after the work tonight, with oplevs centered, and the OSEM sensor values.

Attachment 1: Screenshot_2022-05-20_20-38-25.png
Screenshot_2022-05-20_20-38-25.png
Attachment 2: OSEMs.txt
2022-05-20 20:39:15.780409
 =====BS=====
C1:SUS-BS_ULSEN_OUT16 =   599 
C1:SUS-BS_LLSEN_OUT16 =   575 
C1:SUS-BS_URSEN_OUT16 =   602 
C1:SUS-BS_LRSEN_OUT16 =   636 
C1:SUS-BS_SDSEN_OUT16 =   669 
 =====ITMX=====
C1:SUS-ITMX_ULSEN_OUT16 =   403 
C1:SUS-ITMX_LLSEN_OUT16 =   609 
... 76 more lines ...
  16867   Fri May 20 12:42:23 2022 TegaUpdateVACDoor installation on the end stations

[JC, Tega, Chub]

Today we installed the 200 lbs doors on the end station chambers.

  16866   Thu May 19 19:05:59 2022 AnchalUpdateBHDBS Chamber all work finished, BHD path setup

[Anchal, Paco, Yuta]

BS Oplev Path

  • The changed position of PRM (40m/16863) meant that BS oplev path is getting clipped by the PRM SOS tower.
  • We had to move BSOL ~ 16 cm North and ~ 1.7 cm East.
  • This means that the BS Oplev input beam is now coming behind TT2 instead of infront of it.
  • We also had to align the beam such that input and returning beam are colinear.
  • This meant we, had to change the mount of the upstream beamsplitter in the in-air table so that we can use that for separating the return beam.
  • Again, we should order 2 inc visible BS for this path.
  • Half of the return beam is making its way all the way back to the laser head. I'm not sure if that can be an issue for our oplev loops.
  • We kept the SRM Oplev path same using irises on the table.

PRM Oplev

  • Again, due to changed position of PRM and BS Oplev, it became very hard to setup oplev for PRM.
  • We found a special position which allows us to catch returning beam through the center of the window.
    • But this returning beam is not prompt reflection from PRM, it is reflection of the HR surface.
    • We are hitting about ~5 mm from the edge of PRMOL mirror (because we cannot move the mirror anymore south to avoid clipping BS and SRM input oplev beams)
    • We put in a 1.1m focal length lens in the input beam to narrow the beam on PRMOL so that it doesn't clip
  • We did not put any lens for the return beam. The sensitivity of this oplev might be low due to slighlty bigger beam on the QPD than others (SRM, BS). We can revisit and insert a lens later if required.

Interferometer alignment and PRM alignment

  • The work on BS table did not change the table balance much. We got back the alignment pretty much instantly.
  • We were able to maximize the arm transmissions.
  • Then we used a beam card with hole to check for reflection from PRM and used PRM (mostly pitch correction) to get the return beam back in same way.
  • This recovered REFL beam on the camera. We used REFLDC signal to align PRM better and maximized it.
  • We centered BS, SRM and PRM oplevs after this point.

LO beam mode correction and spatial overlap

  • We tried changing the distance between LO3 and LO4 to get a better output LO beam.
  • We also tried to swap the LO4 mirror with the spare mirror but we had the same result.
  • Eventually, we decided to move LO3 back to East and LO4 to the west edge of the table. This made the beam sizes comparable.
    • Future exercise: We think that LO1 or LO2 might be significantly off-spec in their ROCs which might cause this issue.
    • We should rerun the calculations with the ROC values of LO1 and LO2 written in the datasheets and figure out the correct LO3-LO4 length required.
    • We can make this change in the next vent if required.
  • After the beam sizes were looking approximately similar but more iterations of changing length and realigning are required.

Remaining tasks before pumpdown

  • Push/pull too bright/dark OSEMs in the SOSs (40m/16865).
  • Finish LO beam mode correction and spatial overlap.
  • Center all oplevs, note all beam positions on camera, and note down all DC PD values at proper alignment.
  16865   Thu May 19 18:56:08 2022 yutaUpdateBHDCurrent OSEM sensor values with all the suspensions aligned

Current OSEM sensor values with all the suspensions aligned are attached.
For 'BS','ITMX','ETMX','ITMY','ETMY','PRM','SRM','LO1','LO2', the ones out of the range [200,800] are marked, and for 'PR2','PR3','SR2','AS1','AS4', the ones out of the range [6000,24000] are marked.

Attachment 1: OSEMS.txt
 =====BS=====
C1:SUS-BS_ULSEN_OUT16 =   602 
C1:SUS-BS_LLSEN_OUT16 =   578 
C1:SUS-BS_URSEN_OUT16 =   606 
C1:SUS-BS_LRSEN_OUT16 =   639 
C1:SUS-BS_SDSEN_OUT16 =   672 
 =====ITMX=====
C1:SUS-ITMX_ULSEN_OUT16 =   403 
C1:SUS-ITMX_LLSEN_OUT16 =   606 
C1:SUS-ITMX_URSEN_OUT16 =   679 
... 75 more lines ...
  16864   Thu May 19 08:51:40 2022 PacoUpdateBHDSRM OpLev

[Paco, Ian]

After agreement from Yuta/Anchal, I moved POYM1 yaw to clear the aforementioned path, and Ian restored the POY11 RFPD path. The demodulation phase might need to be corrected afterwards, before any lockign attempts.

Quote:

[Paco]

The SRM Oplev injection and detection paths interfere heavily with the POY11. Due to the limited optical access, I suggest we try steering POYM1 YAW and adapting the RFPD path accordingly.

 

  16863   Wed May 18 17:23:15 2022 AnchalUpdateBHDPlaced PRM in BS Chamber

[Anchal, Paco, Yuta, JC]

SRM Oplev setup

  • We setup SRM oplev path for the aligned position of SRM.
  • This was bit hard, because the return beam was following almost the same path as the input beam, and the return beam had become about 1 cm in diameter.
  • We replaced one of the in-air steering mirror of SRM op-lev input beam with a 1 inch BS on a non-steeerable mount.
  • The returning oplev beam is picked at transmission from this BS.
  • Note: we are not sure if this BS is actually coated for IR or Visible. We couldn't find a visible BS in the lab. We should order a 2 in diameter visible BS to be placed in this position.
  • Half of the input beam would be used for PRM Oplev input.
  • The returning beam was focused with a 100mm focal length lens. Again, this lens is not verified to be for visible wavelength. We think it might have an AR coating for IR. We should get a visible lens for this position also.

PRM Placement

  • PRM placed in nominal position + 2 cm, East.
  • Currently, PRM SOS tower is blocking BS oplev input beam, this needs to be adjusted.
  • Installed PRMOL at nominal position + 2 cm East (to clear path from TT2)
  • I balanced the table succesfully, first using spirit bubble level and then OSEM levels of BS, SR2, PR3 and LO2.
    • Note, that we need to adjust OSEM positions in many of these SOS before pumping down.
  • Input beam from TT2 is going through center of PRM but the reflction is not coming back from PR2, maybe it is missing PR2 or PR2 alignment needs to be adjusted.
  16862   Wed May 18 09:02:52 2022 AnchalUpdateBHDWFS1 PD centered

I centered WFS1 PD so that IMC WFS Servo does not go out of range.

 

  16861   Wed May 18 08:30:29 2022 PacoUpdateBHDSRM OpLev

[Paco]

The SRM Oplev injection and detection paths interfere heavily with the POY11. Due to the limited optical access, I suggest we try steering POYM1 YAW and adapting the RFPD path accordingly.

  16860   Tue May 17 18:43:38 2022 AnchalUpdateBHDPlaced SRM in ITMY Chamber

[Anchal, Paco, Yuta]

SRM Placement

  • SRM was moved from its parked location to the nominal position in the ITMY chamber.
  • This imbalanced the table a lot as all SOS towers ended up on the south side of the table.
  • I needed additionally three SOS tower side walls to recover the balance of the table.
    • I initially tried to use a level meter on my phone which claimed to have 0.1 degrees of accuracy. But it turned out to be a bad idea.
    • Eventually, I used the spirit bubble level meter we have, along with the OSEM values of ITMY, AS1, and AS4.
  • At the end, the table is balanced as it was before, all SOS are damping usually.

SRM Sat Amp Box setup

  • SRM Gold Box Sat Amp was found near the BS chamber.
  • This box was moved to the ITMY chamber.
  • The new flange on the East end was marked earlier for SRM. This flange on the vacuum side was connected with new in-vacuum blue ribbon cables.
  • We had previously moved the cable post for SRM (40m/16849) behind AS2. This cable post is connected to the old in-vacuum cable.
  • It would have changed the table balance to remove this cable post and connect new in-vacuum cables to it, so we decided to do this in the next vent when we put the BHD board on the table.
  • For now, we connected the old in-vacuum cable to the new in-vacuum blue ribbon cables inside.
    • Note, that the old in-vacuum cable has a gender flipping section which also mirrors the pin layout.
    • We installed pin mirroring cables on the outside between the Sat Amp Box and the vacuum flange to revert back the additional mirroring.
    • However it happened, now the Sat Amp Box is working, with all OSEMs and coils alive.
  • One peculiarity we found was that the SRM face OSEMs have only about 250-300 um of range, which is roughly 3 times less than the other OSEMs in other SOSs.
  • SRM side OSEM however behaves normally.
  • After installment, at the free-hanging state, SRM LL OSEM is saturated (too bright) and other face OSEMs are close to total brightness state.
  • We'll first put the alignment offsets to get the SRM perpendicular to the beam coming from SR2 and then center the OSEMs in this tiny range.
  • The low OSEM range could be due to improper biasing from the Sat Amp Box. Hopefully, with new electronics, this issue would go away in future.
  16859   Mon May 16 19:14:17 2022 AnchalUpdateBHDCamera set on AS path and BHDBS output path

[Anchal, Paco, Yuta]

  • We aligned AS path avoiding any clipping to the AP table where we setup a camera with a lens.
    • To do this we had to move AS6 in North direction for ~1cm.
    • The Injection table was imbalanced by this move to drop the IMC transmission to half.
    • We did not balance the table again, we steered the input mirror to reach to 1000 counts (out of 1200 nominal) and then used WFS loop to get to the last bit.
    • The input to the arm cavities did not change much, XARM was still flashing to 0.8 max height and YARM to 0.2. We recovered these easily using the cavity mirror pair.
  • We aligned the LO beam to be spatially matched on BHDBS with AS path.
    • The LO beam was steered to roughly overlap with the AS beam outputs on the BHDBS.
    • However, the LO beam size is very large and diverges after LO4.
    • According to 40m/15379, the 0.15m ROC of LO4 right after the beam waist is supposed to collimate the beam to a 522 um waist.
    • We confirmed that LO4 is marked as a 0.15m ROC mirror on its edge and the HR coating is facing the incident beam.
      • Conjecture (AG): The coating was applied to the flat side of the optic instead of the curved side.
      • This would explain why the beam is continuing to diverge after reflecting from LO4, and diverging fast.
    • We need to fix this issue before pumping down otherwise the mode matching would be too poor in BHDBS to have any meaningful results.
  • The output of BHDBS was steered out and a GigE camera is set up to see this path.
    • The camera is set to see the transmitted AS beam from BHD BS (and reflected LO beam).
    • But the camera is unable to see any LO beam due to large divergence.
    • The LO beam essentially disappears after ~30 cm from the BHDBS.

 

  16858   Mon May 16 16:13:01 2022 YehonathanUpdateBHDInitial BHD modeling: Damped suspension model

I was finally able to set up a stable suspension model with the help of Yuta and I'm now ready to start doing some MICH noise budgeting with BHD readout. (Tip: turns out that in the zpk function in Matlab you should multiply the poles and zeros by -2*pi to match the zpk TFs in Foton)

I copied all the filters from the suspension MEDM screens into a Matlab. Those filters were concatenated with a single pendulum suspension TF with poles at [0.05e-1+1i, 0.05e-1-1i] and a gain of 4 N/kg.

I multiplied the OLTF with the real gains at the DAC/DAC/OSEMs/Coil Driver and Coils. I ignore whitening/dewhitening for now. The OLTF was calculated with no additional ad-hoc gain.

Attachment 1 shows the calculated open-loop transfer function.

Attachment 2 shows OLTF of ETMY measured last week.

Attachment 3 shows the step and impulse responses of the closed-loop system.

 

 

Attachment 1: Damped_SUS.png
Damped_SUS.png
Attachment 2: ETMY_SUSPOS_GOL.pdf
ETMY_SUSPOS_GOL.pdf ETMY_SUSPOS_GOL.pdf
Attachment 3: SUS_Response.png
SUS_Response.png
  16857   Mon May 16 14:46:35 2022 TommyUpdateElectronicsRFSoC MTS Work

We followed the manual's guide for setting up MTS to sync on external signal. In the xrfdc package, we update the RFdc class to have RunMTS, SysRefEnable, and SysRefDisable functions as prescribed on page 180 of the manual. Then, we attempted to run the new functions in the notebook and read the DAC signal outputs on an oscilloscope. The DACs were not synced. We were also unable to get FIFOlatency readings. 

  16856   Mon May 16 13:22:59 2022 yutaUpdateBHDREFL and AS paths aligned at AP table

After Xarm and Yarm were aligned by Anchal et al, I aligned AS and REFL path in the AP table.
REFL path was alreasy almost perfectly aligned.

REFL path
 -REFL beam centered on the REFL camera
 -Aligned so that REFL55 and REFL33 RFPDs give maximum analog DC outputs when ITMY was misaligned to avoid MICH fringe
 -Aligned so that REFL11 give maximum C1:LSC-REFL11_I_ERR (analog DC output on REFL11 RFPD seemed to be not working)

AS path
 -AS beam centered on the AS camera. AS beam seems to be clipped at right side when you see at the viewport from -Y side.
 -Aligned so that AS55 give maximum C1:LSC-ASDC_OUT16 (analog DC output on AS55 RFPD seemed to be not working)
 -Aligned so that AS110 give maximum analog DC output

Attachment 1: REFLPOP.JPG
REFLPOP.JPG
Attachment 2: POPAS.JPG
POPAS.JPG
  16855   Mon May 16 12:59:27 2022 ChrisUpdateDAQDAQ troubleshooting

It looks like the RFM problem started a little after 2am on Saturday morning (attachment 1). It’s subsequent to what I did, but during a time of no apparent activity, either by me or others.

The pattern of errors on c1rfm (attachment 2) looks very much like this one previously reported by Gautam (errors on all IRFM0 ipcs). Maybe the fix described in Koji’s followup will work again (involving hard reboots).

Attachment 1: timeseries.png
timeseries.png
Attachment 2: err.png
err.png
  16854   Mon May 16 10:49:01 2022 AnchalUpdateDAQDAQ troubleshooting

[Anchal, Paco, JC]

Thanks Chris for the fix. We are able to access the testpoints now but we started facing another issue this morning, not sure how it is related to what you did.

  • The C1:LSC-TRX_OUT and C1:LSC-TRY_OUT channels are stuck to zero value.
  • These were the channels we used until last friday to align the interferometer.
  • These channels are routed through the c1rfm FE model (Reflected Memory model is the name, I think). These channels carry the IR transmission photodiode monitors at the two ends of the interferometer, where they are first logged into the local FEs as C1:SUS-ETMX_TRX and C1:SUS-ETMY_TRY .
  • These channels are then fed to C1:SCX-RFM_TRX -> C1:RFM_TRX -> C1:RFM-LSC_TRX -> C1:LSC-TRX and similar for Y side.
  • We are able to see channels in the end FE filtermodule testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT)
  • However, we are unable to see the same signal in c1rfm filter module testpoints like C1:RFM_TRX_IN1, C1:RFM_TRY_IN1 etc
  • There is an IPC error shown in CDS FE status screen for c1rfm in c1sus. But we remember seeing this red for a long time and have been ignoring it so far as everything was working regardless.

The steps we have tried to fix this are:

  • Restart all the FE models in c1lsc, c1sus, and c1ioo (without restarting the computers themselves) , and then burt restore.
  • Restart all the FE models in c1iscex, and c1iscey (only c1iscey computer was restarted) , and then burt restore.

These above steps did not fix the issue. Since we have  the testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT) for now to monitor the transmission levels, we are going ahead with our upgrade work without resovling this issue. Please let us know if you have any insights.

  16853   Sat May 14 08:36:03 2022 ChrisUpdateDAQDAQ troubleshooting

I heard a rumor about a DAQ problem at the 40m.

To investigate, I tried retrieving data from some channels under C1:SUS-AS1 on the c1sus2 front end. DQ channels worked fine, testpoint channels did not. This pointed to an issue involving the communication with awgtpman. However, AWG excitations did work. So the issue seemed to be specific to the communication between daqd and awgtpman.

daqd logs were complaining of an error in the tpRequest function: error code -3/couldn't create test point handle. (Confusingly, part of the error message was buffered somewhere, and would only print after a subsequent connection to daqd was made.) This message signifies some kind of failure in setting up the RPC connection to awgtpman. A further error string is available from the system to explain the cause of the failure, but daqd does not provide it. So we have to guess...

One of the reasons an RPC connection can fail is if the server name cannot be resolved. Indeed, address lookup for c1sus2 from fb1 was broken:

$ host c1sus2
Host c1sus2 not found: 3(NXDOMAIN)

In /etc/resolv.conf on fb1 there was the following line:

search martian.113.168.192.in-addr.arpa

Changing this to search martian got address lookup on fb1 working:

$ host c1sus2
c1sus2.martian has address 192.168.113.87

But testpoints still could not be retrieved from c1sus2, even after a daqd restart.

In /etc/hosts on fb1 I found the following:

192.168.113.92  c1sus2

Changing the hardcoded address to the value returned by the nameserver (192.168.113.87) fixed the problem.

It might be even better to remove the hardcoded addresses of front ends from the hosts file, letting DNS function as the sole source of truth. But a full system restart should be performed after such a change, to ensure nothing else is broken by it. I leave that for another time.

  16852   Fri May 13 18:42:13 2022 PacoUpdateAlignmentITMX and ITMY sat amp failures

[Yuta, Anchal, Paco]

As described briefly by JC, there were multiple failure modes going during this work segment. blush


ITMX SatAmp SAGA

Indeed, the 64 pin crimp cable from the gold sat amp box broke when work around ITMX chamber was ongoing. We found the right 64 pin head replacement around and moved on to fix the connector in-situ. After a first attempt, we suddenly lost all damping on vertex SUS (driven by these old sat amp electronics) because our c1susaux acromag chassis stopped working. After looking around the 1x5 rack electronics we noted that one of the +- 20 VDC Sorensens were at 11.6 VDC, drawing 6.7 A of current (nominally this supply draws over 5 Amps!) so we realized we had not connected the ITMX sat amp correctly, and the DC rail voltage drop busted the acromag power as well, tripping all the other watchdogs devil ...

We fixed this by first, unplugging the shorted cable from the rack (at which point the supply went back to 20 VDC, 4.7 A) and then carefully redoing the crimp connector. The second attempt was successful and we restored the c1susaux modbusIOC service (i.e. slow controls).


ITMY SatAmp SAGA

As we restored the slow controls, and damped most vertex suspensions, we noticed ITMY UL and SD osems were reading 0 counts both on the slow and fast ADCs. crying We suspected we had pulled some wires around when busy with the ITMX sat amp saga. We found that Side OSEM cLEMO cable was very loose on the whitening board. In fact, we have had no side osem signal on ITMY for some time. We fixed this. Nevertheless the UL channel remained silent... We then did the following tests:

  • Test PD mon outputs on the whitening card. We realized the whitening cards were mislabeled, with ITMX and ITMY flipped angry. We have labeled them appropriately.
  • Tested input DB15 cable with breakout board.
  • Went to the ITMY sat amp box and used the satellite box TESTER 2 on J1. It seemed correct.
  • We opened the chamber, tested the in-vacuum segments, they all were ok.
  • We flipped UR-UL OSEMs and found that the UL OSEM is healthy and works fine on UR channel.
  • We tested the in-air cable between satellite box and vacuum flange and it was ok too.
  • We suspected that the satellite box tester  is lying, so we replaced the satellite box with the spare old MC1 satellite box, and indeed that solved the issue.

DO NOT TRUST THE SATELLITE BOX TESTER 2.


Current state:

  • IMC locking normally.
  • All suspensions are damping properly.
  • Oplevs are not centered.
  • No flashing on either of the arms. We had no luck in ~20 min of attempt with just input injection changed.
  • On kicking PR3, we do see some flashing on XARM, which means XARM cavity atleast is somewhat aligned.
  • All remaining tasks before pumpdown are still remaining. We just lost the whole day.
  16851   Fri May 13 14:26:00 2022 JCUpdateAlignmentLO2 Beam

[Yehonathan, JC]

Yehonathan and I attempted to align the LO2 beam today through the BS chamber and ITMX Chamber. We found the LO2 beam was blocked by the POKM1 Mirror. During this attempt, I tapped TT2 with the Laser Card. This caused the mirror to shake and dampen into a new postion. Afterwards, when putting the door back on ITMX, one of the older cables were pulled and the insulation was torn. This caused some major issues and we have been able to regain either of the arms to their original standings.

  16850   Thu May 12 20:24:29 2022 yutaUpdateBHDPOX and POY investigation

[Anchal, Yuta]

We checked POX and POY RF signal chains for sanity check since Xarm cannot be locked in IR stably as opposed to Yarm.
POX beam seems to be healthy. This issue doesn't prevent us from closing the vacuum tank.

POY
 - RF PD has SPB-10.7+ and ZFL-500NL+ attached to the RF output.
 - At the demodulation electronics rack, SMA connectors are used everywhere.
 - With Yarm flashing at ~1, RF output has ~24 mVpp right after RF PD, ~580mVpp after SPB-10.7+ and ZFL-500NL+, and ~150mVpp at right before the demodulation box.
 - There is roughly a factor of 3 loss in the cabling from POY RF PD to the demodulation rack.
 - Laser power at POY RF PD was measured to be 16 uW

POX
 - RF PD doesn't have amplifiers attached.
 - At the demodulation electronics rack, N connector is used.
 - With Xarm flashing at ~1, RF output has ~30 mVpp right after RF PD, and ~20mVpp at right before the demodulation box.
 - Losses in the cabling from POX RF PD to the demodulation rack is small compared with that for POY.
 - Laser power at POX RF PD was measured to be 16 uW

Summary
 - POX and POY RF PDs are receiving almost the same mount of power
 - POY has larger error signal than POX because of RF amplifier, but the cable loss is high

Conclusion
 - There might be something in the electronics, but we can close the vacuum tanks

Attachment 1: POY.JPG
POY.JPG
  16849   Thu May 12 20:11:18 2022 AnchalUpdateBHDBHDBS Output beams steered out to ITMY table

I successfully steered out the two output beams from BHD BS to ITMY table today. This required significant changes on the table, but I was able to bring back the table to balance coarsely and then recover YARM flashing with fine tuning of ITMY.

  • The counterweights were kept at the North end of the table which was in way of one of the output beams of BHD.
  • So I saved the level meter positions in my head and removed those counterweights.
  • I also needed to remove the cable post for ITMY and SRM that was in the center of the table.
  • I installed a new cable post which is just for SRM and is behind AS2. ITMY's cable post is next to it on the other edge of the table. This is to ensure that BHD board can come in later without disturbing existing layout.
  • I got 3 Y1-45P and 1 Y1-0 mirror. The Y1-0 mirror was not installed on a mount, so I removed an older optic which was unlabeled and put this on it's mount.
  • Note that I noticed that some light (significant enough to be visible on my card) is leaking out of the 45P mirrors. We need to make sure we aren't loosing too much power due to this.
  • Both beams are steered through the center of the window, they are separating outside and not clipping on any of the existing optics outside. (See attachment 1, the red beam in the center is the ITMY oplev input beam and the two IR beams are the outputs from BHD BS).
  • Also note that I didn't find any LO beam while doing this work. I only used AS beam to align the path.
  • I centered the ITMY oplev at the end.

Next steps:

  • LO path needs to be tuned up and cleared off again. We need to match the beams on BHD BS as well.
  • Setup steering mirrors and photodiodes on the outside table on ITMY.
Attachment 1: signal-2022-05-12-201844.jpeg
signal-2022-05-12-201844.jpeg
  16848   Thu May 12 19:55:01 2022 TegaUpdateBHDAS path alignment

[Yuta, Tega]

We finally managed to steer the AS beam from ITMY chamber, through BS and IMC chambers, to the in-air AP table.

We moved the AS5 mirror north to its nominal position and we also moved the ASL lens on BS chamber back to its nominal position. Attached photos are taken after today's alignment work.

Attachment 1: AS.JPG
AS.JPG
Attachment 2: OMCchamber.JPG
OMCchamber.JPG
  16847   Thu May 12 17:20:08 2022 PacoUpdateAlignmentPOP Beam on CCD

[Paco]

Got POP beam centered on camera and nominally on the two PDs. Attachment #1 shows "carrier" camera.

Attachment 1: PXL_20220513_002047233.jpg
PXL_20220513_002047233.jpg
  16846   Thu May 12 13:46:59 2022 JcUpdateAlignmentPOP Beam

[Tega, JC]

Tega and I went in to adjust the POP being in the ITMX Table. The beam entered the table high, so we adjusted the this by adding mirrors (The highlighted in Turqoise are mirrors which adjust the pitch of the beam). All the mirrors are set and we are now in the process of adjusting the PD.

Attachment 1: IMG_0777.jpeg
IMG_0777.jpeg
  16845   Wed May 11 15:49:42 2022 JCUpdateOPLEV TablesGreen Beam OPLEV Alignment

[Paco, JC]

Paco and I began aligning the Green Beam in the BS Oplev Table. while aligning the GRN-TRX, the initial beam was entering the table a bit low. To fix this, Paco went into the chamber and correcting the pitch with the steering mirror. The GRN-TRX is now set up, both the PD and Camera. Paco is continuing to work on the GRN-TRY and will update later on today. 

In the morning, I will update this post with photos of the new arrangement of the BS OPLEV Table.


Update Wed May 11 16:54:49 2022

[Paco]

GRY is now better mode matched to the YARM and is on the edge of locking, but it more work is needed to improve the alignment. The key difference this time with respect to previous attempts was to scan the two lenses on translation stages along the green injection path. This improved the GTRY level by a factor of 2.5, and I know it can be further improved. Anyways, the locked HOMs are nicely centered on the GTRY PD, so we are likely done with the in-vac GTRY GTRX alignment.


Update Wed May 12 10:59:22 2022

[JC]

The GTRX PD is now set up and connected. The camera have been set to an angle because the cable to connect it is too thick for the camera to maintain its original position along the side. 

 

Attachment 1: IMG_0770.jpeg
IMG_0770.jpeg
  16844   Tue May 10 18:25:43 2022 PacoUpdateBHDGreen Transmission path

[Yuta, Paco]

We installed GRX_SM1, GRX_SM2, and finished aligning the GRY_SM1, and GRY_SM2 steering mirrors in the BS and IMC Chambers. GRY_SM1 was slightly misplaced (by ~ 2 inches), so we had to move it slightly. Luckily this didn't grossly misaligned the IMC, and we could recover quickly by touching MC1 & MC3 pitch, and MC1 slight yaw.

Then, Yuta installed GRX_SM1, and GRX_SM2 by repurposing two 45 AOI P-Pol GR transmission mirrors on the flowbench. Because one of the weights on the BSC was in the way of GRX_SM2, it was shifted it before installation. This probably shifted the balancing of the whole table. The GRY beam is still not lock-able to the YARM, so as a proxy for GRY transmission beam we used the slight GRX reflection from the BS, and noted slight clipping through PR3 (in transmission). This should probably be checked with GTRY.

We believe this is the last installation operation of this vent.


We made sure the WFS feedback loop is working, and realigned the arm cavities to be flashing.

  16843   Tue May 10 16:31:14 2022 TegaUpdateBHDITMX optlev return beam steered to QPD

Followed the steps below to complete the ITMX optlev installation. The ITMX optlev return beam now reaches its QPD without being blocked by the input steering mirror.

Although, I centered the ITMX optlev readout, this was not done when the XARM flashing is maximized bcos the IMC chamber was being worked on, so this should be done later when the IR beam is back.

Quote:

Next steps:

  • The Oplev beam paths need to be adjusted.
    • The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
    • First setup two irises to save ingoing path.
    • Then make space for the returning beam by changing the steering mirror positions.
    • Then recover the ingoing path to the center of irises.
    • Steer the returning beam to the QPD.
    • Then maximize the flashing on XARM and center the oplev to save this position.
  16842   Tue May 10 15:46:38 2022 JCUpdateBHDRelocate green TRX and TRY components from PSL table to BS table

[JC, Tega]

Tega and I cleaned up the BS OPLEV Table and took out a couple of mirrors and an extra PD. The PD which was removed is "IP-POS - X/Y  Reversed". In addition to this, the cable is zip-tied to the others located on the outside of the table in case this is required later on. 

Next, we placed the cameras and mirrors for the green beam into their postions. A beam splitter and 4 mirrors were relocated from PSL table and placed onto the BS Oplev table to complete this. I will upload the picture of the newly updated photo with arrows of the beam routes.

Attachment 1: IMG_0741.jpeg
IMG_0741.jpeg
Attachment 2: IMG_0753.jpeg
IMG_0753.jpeg
  16840   Mon May 9 23:18:44 2022 KojiUpdateBHDBHD Platform Progress status

I quickly ran the FEA model to check the resonant freqs of the BHD platform.
The boundary conditions were:

  • The platform was not loaded
  • FIxed constraints were given to the five legs

Don has optimized the cut-out size for the OMCs to increase the rigidity of the plate. Also, the ribbed grid is made at the bottom side.

The lowest mode is at 168Hz. Because there is no leg around, it seems reasonable to have this kind of mode as the fundamental mode.
The other mode lined up at 291Hz, 394Hz, 402Hz, ...
The mode freqs will be lower once the platform is loaded. But as the unloaded platform mode, these mode freqs sound pretty good numbers.

Attachment 1: Screenshot_2022-05-09_224145.png
Screenshot_2022-05-09_224145.png
Attachment 2: Screenshot_2022-05-09_224137.png
Screenshot_2022-05-09_224137.png
Attachment 3: bhd_platform_168.png
bhd_platform_168.png
Attachment 4: bhd_platform_291.png
bhd_platform_291.png
Attachment 5: bhd_platform_394.png
bhd_platform_394.png
Attachment 6: bhd_platform_402.png
bhd_platform_402.png
  16839   Mon May 9 22:19:06 2022 KojiUpdateBHDBHD Platform Progress status

[Don, Koji]

Don is working on finalizing the BHD Platform design. All the components on the BHD platform are almost populated and aligned.

Don is still working on the table legs so that we can detach the legs when we need to float the table in the future.
The BHD BS mount will have a third picomotor so that we can steer 3 dof with the mount while the remaining dof needs to be provided by the OMC.
The BHD BS position is going to be adjusted so that the incident and trans beams have sufficient clearance.
The OMC legs (kinematic mounts) need more work so that we can adjust their positions for initial setup while they can be the reference for the reproducible placement of the OMCs.
The OMCs are rigidly held with the legs. For the damping of the 1-kHz body bode, which has a relatively high Q, there will be a dissipative element touching the glass breadboard.

Attachment 1: Screenshot_2022-05-09_220220.png
Screenshot_2022-05-09_220220.png
Attachment 2: Screenshot_2022-05-09_222405.png
Screenshot_2022-05-09_222405.png
  16838   Mon May 9 18:49:05 2022 TegaUpdateBHDRelocate green TRX and TRY PDs/cams/optics from PSL table to BS table

[Paco, Tega]

Started work on the relocating the green transmission optics, cameras and PDs. Before removing the any of the optics, we checked and confirmed that the PDs and Cams are indeed connected to the GRN TRX/Y medm channels. Then added labels to the cables before moving them.

Plumbing:

  • Moved all power and signal cables for the PDs and cameras from PSL table to BS table. See attachment #1

Relocated Optics & PDs & Cameras:

  • TRX and TRY cameras
  • TRX and TRY PDs
  • 1 BS, 2 lens for PDs and a steering mirror, see Atachement #2

 

Attachment 1: IMG_20220509_184439943.jpg
IMG_20220509_184439943.jpg
Attachment 2: IMG_20220509_184154722.jpg
IMG_20220509_184154722.jpg
  16837   Mon May 9 18:43:03 2022 AnchalUpdateBHDITMX table layout corrected

As I went to correct the ITMX Oplev mirrors, I found that both mirrors were placed in very different positions than the design position. Part of the reason I think was to preserve outside oplev path, and party because a counterweight was in ITMXOL1 position. I had to do following steps to correct this:

  • I noted down level meter readings of the table before making any changes.
  • I removed the counter weight from near the center of the table.
  • I placed the Oplev mirrors in the nominal positions.
  • I placed the counter weight near previous position.
  • I moved a edge hanging counter weight to get back the level meter to its previous state coarsely.
  • Then I used dataviewer to find the previous OSEM PD monitor values and changed ITMX PIT and YAW to come closer to those PD values. And voila, I regained the flashing on Xarm. I nudged the ITMX pit and yaw bit more to maximize it.
  • I then went back to aligning the Oplevs properly.
  • Then I adjusted the POP mirrors to get the beam back through center of window. This was very tricky and took a lot of time.
  • Now the beam is going through near center and the oplev beams are far away enough from POP_SM5.
  • On the outside table, I noted the POP beam and the oplev beam. I corrected the pit of the returning beam to get the oplev beam at nominal height on outside table.

ITMX Sat Amp is flaky

[Anchal, Paco]

During the above work, i must have kicked the cable between the vacuum flange and the satellite amplifier box for ITMX. This disconnected all the OSEMs and Coils. We tried several things to debug this and finally found that nudging the connections on Sat Amp box brought the OSEMs and coils back online. Note that the connector was not partially out or in a state that obviously showed disconnection of the pins. I'm glad we are putting in new electronics soon for the vertex optics as well.


Next steps:

  • I showed Tega the returning oplev beam and the POP beam coming out of the ITMX chamber.
  • The Oplev beam paths need to be adjusted.
    • The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
    • First setup two irises to save ingoing path.
    • Then make space for the returning beam by changing the steering mirror positions.
    • Then recover the ingoing path to the center of irises.
    • Steer the returning beam to the QPD.
    • Then maximize the flashing on XARM and center the oplev to save this position.
  • POP beam needs to be directed to previous setup on far side of table.
    • The POP beam is coming out at the rising angle.
    • This is good for us if we do bit unconventional stuff and transfer the beam to other side of table at an elevated height. Given how close all the beams are coming out of the viewport, I think this is the best solution in terms of saving time.
    • Get the beam down to the old setup which was camera and photodiodes all aligned.

 

 

ELOG V3.1.3-